OpenStudy (anonymous):

1/2x+1/4x=16 16 21.3 32 64

OpenStudy (anonymous):

Hello, try multiplying the whole equation by 4x (the largest demoniator) and then solve for x

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

And what did you get?

OpenStudy (anonymous):

i didnt get the answer

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

Or where did you get stuck?

OpenStudy (anonymous):

the first one i got 12

OpenStudy (anonymous):

the second one i got 15.975

OpenStudy (anonymous):

the third one i got 24

OpenStudy (anonymous):

wait I think I read the problem wrong: is it (1/2) x + (1/4)x = 16 or is it 1/(2x) + 1/(4x) = 16... note there is a big difference

OpenStudy (anonymous):

If it is the first case (as I now suspect) try multiplying the whole equation by 4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

So when you multiply the equation by 4 do you get: \[4( \frac{1}{2}x + \frac{1}{4} x = 16) \ \ \ \ \rightarrow \ \ \ \ 2x+x=16*4\]